OpenStudy (anonymous):

Given the function g(x) = -3x + 4, compare and contrast g(-2) and g(4). Choose the statement that is true concerning these two values. Answer A:The values of g(-2) and g(4) cannot be compared. B:The value of g(-2) is the same as the value of g(4). C:The value of g(-2) is larger than the value of g(4). D:The value of g(-2) is smaller than the value of g(4).

sam (.sam.):

C

OpenStudy (anonymous):

how did u get it

sam (.sam.):

g(x) = -3x + 4 -------------------------------------------------------- g(-2) = -3(-2) + 4=10 g(4) = -3(4) + 4=-8

sam (.sam.):

g(-2) > g(4)

sam (.sam.):

10 > -8

OpenStudy (anonymous):

If f(x) = 2x + 8, find f(9).

OpenStudy (anonymous):

can you do thiss too plzz